Solve for the determinant in the equation below

2 −2 1

2 2 1

4 1 3


1
Expert's answer
2022-03-16T17:56:06-0400

Solution

"det(A)=\begin{vmatrix}
 2 & -2 & 1\\
 2 & 2 & 1\\
 4 & 1 & 3\\
\end{vmatrix}"


"det(A)=2\begin{vmatrix}
 2 & 1 \\
 1 & 3
\end{vmatrix}-(-2)\begin{vmatrix}
 2 &1 \\
 4 & 3
\end{vmatrix}+(1)\begin{vmatrix}
 2 & 2 \\
 4 & 1
\end{vmatrix}"


"det(A)=(2)(6-1)+(2)(6-4)+(1)(2-8)"


"det(A)=(2)(5)+(2)(2)+(1)(-6)"


"det(A)=10+4-6"


"det(A)=8"


Hence


"\begin{vmatrix}
 2 & -2 & 1\\
 2 & 2 & 1\\
 4 & 1 & 3\\
\end{vmatrix}=8"